if F(x) = 7/(x-4) and g(x) = 2/x find the composite of f of g (x) and the domain
Answer by edjones(8007)

f(x) = 7/(x-4) and g(x) = 2/x
f(g(x))=7/(2/x - 4)
=7/((2-4x)/x)
=7 * x/(2-4x)
=7x/(2-4x)
.
domain: all real numbers except x=1/2
.
Ed